OpenStudy (help!!!!):

a danish krone is worth 50 french francs; an egyptian pound is worth 17 1/2 krone; there are 350 french francs to $1. How many egyptian pounds are equivalent to $1?

OpenStudy (help!!!!):

How would I set up the equation?

OpenStudy (anonymous):

Set up all of the equations first: K=50F, F=K/50 E=17.5K, K=E/17.5 350F=D \[D=350F=350(\dfrac{K}{50})=350(\dfrac{(\dfrac{E}{17.5})}{50})=\dfrac{7}{17.5}E\]

OpenStudy (help!!!!):

nvm I got it! Thanks for the help!!